OpenStudy (anonymous):

solve for k the system kx-5y=8 7x+5y=10 is consistent.

OpenStudy (anonymous):

kx - 5y = 8 7x + 5y = 10 -------------- Try adding the equations together and see what you get

OpenStudy (anonymous):

mark_o, please do not give the answer to the asker, but help them receive the answer by themselves as stated in the Code of Conduct.

OpenStudy (anonymous):

ah ok i thought you are talking to me sorry chandhooguy

OpenStudy (anonymous):

That's ok

OpenStudy (anonymous):

bec, i myself cant figure what k is,,lol

OpenStudy (anonymous):

;) I can help. Would you like me to?

OpenStudy (anonymous):

yes, plz do it for us

OpenStudy (anonymous):

kx - 5y = 8 7x + 5y = 10 -------------- If I add both equations to each other, I get... kx + 7x = 18 Which can be simplified to... (k + 7)x = 18 Helpful?

OpenStudy (anonymous):

a lot better now ,so k will be a function of x right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

ok you got it....it very clear now thnx for your help..lol

OpenStudy (anonymous):

No prob!